A bag contains 3 red, 7 black, 5 green marbles. Four marbles are selected without replacement. What is the probability that all 4 marbles are black?


1
Expert's answer
2021-07-14T17:28:01-0400

"N=3+7+5=15"

"P(B=4)=P(B_1\cap (B_2|B_1)\cap (B_3|(B_1\cap B_2))\cap (B_4|(B_1\cap B_2\cap B_3)))" B-black

"=\frac{7}{15}\times\frac{6}{14}\times\frac{5}{13}\times\frac{4}{12}"

"=\frac{1}{39}"